OpenStudy (anonymous):

f(theta)=sin(theta)...solve f(theta/2)....theta =60 degrees

hero (hero):

\(f(\theta) = \sin(\theta)\) \(f \left(\dfrac{\theta}{2}\right) = \sin \left(\dfrac{\theta}{2} \right)\) \(\theta = 60\) \(f \left(\dfrac{60}{2}\right) = \sin \left(\dfrac{60}{2} \right)\) \(f(30) = \sin(30)\) \(f(30) = \dfrac{1}{2}\)